Consider any real number such that it's square is the same as seventeen less than ten times the number. Find the exact value of the number(s).
The number(s) is(are) =

Answer by jorel555(1290)

n²=10n-17
n²-10n=17
n²-10n+25=42
(n-5)²=42
n=√42+5=11.48 ☺☺☺☺